luhivqqcherry:

helpp withh moree mathh workk . Find the sum: 4 + 4.8 + 5.76 + 6.912 + ... a) Infinity b) 20 c) 24

Midnight97:

This is a sequence where each term after the first one is obtained by multiplying the previous term by 1.2 (because 4 * 1.2 = 4.8, 4.8 * 1.2 = 5.76, and so on)

Midnight97:

For a geometric sequence, the sum S can be found using the formula... S= a/1-r where a is the first term (4) and r is what we multiply it by each time (so 1.2) The finite sum only exists if r < 1 and since our r is 1.2 which is greater than 1...

luhivqqcherry:

my brain is currently in a different world 😭 .

Midnight97:

@luhivqqcherry wrote:
my brain is currently in a different world 😭 .
Lol

Midnight97:

If each term in the sequence keeps getting larger as we multiply by 1.2.... and it never stops

luhivqqcherry:

OHHHHHHHHHHHHHHHHHHHHHHH .

luhivqqcherry:

SO THAT MEANS ITS INFINITY .

Midnight97:

Yes

luhivqqcherry:

MUAJAHAHA TYTYTYYYYY .

Midnight97:

@luhivqqcherry wrote:
MUAJAHAHA TYTYTYYYYY .
You're so welcome (:

luhivqqcherry:

OKAY I HAVE ONE MORE quESTION AND IM DONE .

Midnight97:

@luhivqqcherry wrote:
OKAY I HAVE ONE MORE quESTION AND IM DONE .
Alright let's do it